### v1

@@ Abstract
We present HAHA - a novel approach for animatable human avatar generation from monocular input videos. The proposed method relies on learning the trade-off between the use of Gaussian splatting and a textured mesh for efficient and high fidelity rendering. We demonstrate its efficiency to animate and render full-body human avatars controlled via the SMPL-X parametric model. Our model learns to apply Gaussian splatting only in areas of the SMPL-X mesh where it is necessary, like hair and out-of-mesh clothing. This results in a minimal number of Gaussians being used to represent the full avatar, and reduced rendering artifacts. This allows us to handle the animation of small body parts such as fingers that are traditionally disregarded. We demonstrate the effectiveness of our approach on two open datasets: SnapshotPeople and X-Humans. Our method demonstrates on par reconstruction quality to the state-of-the-art on SnapshotPeople, while using less than a third of Gaussians. HAHA outperforms previous state-of-the-art on novel poses from X-Humans both quantitatively and qualitatively.
Project page: https://david-svitov.github.io/HAHA_project_page/
Figure 1: Optimizing the number of Gaussians. HAHA jointly optimizes a Gaussian splatting model with a textured mesh to improve the photometric quality of the avatars. The method filters out superfluous Gaussians in a learnable, unsupervised manner. As result we can more efficiently and better animate highly articulated parts of a body.
@@ 1 Introduction
To date, several options are available to generate human avatars for what concerns input data. To get the best quality many methods relies on multi-view data [9, 11, 65, 25, 18]. However, complex acquisition systems such as a multi-camera capture setup [23] or a 3D scanner [4] are required to collect such data. On the other hand, some methods [55, 19, 51, 8] use a single image of a person as input, which overcomplicates the task with the necessity to restore unobserved regions of the body. Eventually, the most convenient way is to generate avatars from monocular videos. Using a monocular video provides a trade-off between the complexity of obtaining input data and the quality of the avatar.
@@
  • We develope an unsupervised method for significantly reducing the amount of Gaussians in the scene through the use of textured mesh;
  • We demonstrate that our method can efficiently handle the animation of hands and other highly articulated parts without the need of any additional engineering.
@@ 2 Related work
Several previous works [67, 53] solve the task of generating human avatars with articulated finders using multi-view data. AvatarReX [67] uses a separate parametric model to process hands. As input, they accept a multi-view video of a person. X-Avatar [53] uses a part-specific deformer network to handle the hands. As input X-Avatar gets 3D scans or RGB-D video with depth information. Both methods reconstruct an avatar as a textured mesh that in general leads to more blurried results than 3DGS. In contrast to the previous works, we use only monocular RGB data.
@@
In this work, we research the new task of merging a novel Gaussian-based approach with a classical RGB texture-based. This allows us to reduce the number of Gaussians and therefore reduce memory requirements to store an avatar. Utilizing textured mesh where possible helps us reduce the number of artifacts connected with redundant Gaussians while remaining Gaussians represent out-of-the-mesh elements of avatars. Thus, we leverage the pros from both representations.
@@ 3.2 Monocular avatar training
Figure 3: Stages of training. a) SMPL-X with optimizeble RGB texture fitted on input video frames. b) 3DGS trained as described in Section 3.1. c) All unnecessary Gaussians are deleted (Section 3.3) to merge this step with (a) and get (d).
First stage: Gaussian avatar training. In the first stage (Fig. 2 (a)), we train the 3DGS representation of an avatar by optimizing only local Gaussians transformations $\mu_{i}^{j},r_{i}^{j},s_{i}^{j}$ and color $c_{i}^{j}$. Opacity $o_{i}^{j}$ is fixed to $1$ during this stage as we keep all Gaussians untransparent to efficiently back-propagate image space losses to the SMPL-X parameters. Thus, we force the model to optimize the pose and shape of the underlying mesh rather than deleting Gaussians. We use randomly colored backgrounds in this stage to prevent Gaussians from learning background color.
To optimize Gaussians we use several image space losses as $L_{2}$ loss, $L_{\textrm{LPIPS}}$ perceptual loss [63], $L_{\textrm{SSIM}}$ structure similarity loss, and $L_{\textrm{Sobel}}$ loss to get sharper edges. To calculate $L_{\textrm{Sobel}}$ loss we measure $L_{2}$ between results of applying the Sobel operator [30] to rendered and ground truth images. In other words, we calculate the distance between discrete derivatives of images to account for high-frequency details. We follow [38] and apply $L_{\textrm{KNN}}$, a KNN-based regularization to get smoother results with fewer artifacts. In KNN-regularization we minimize standard deviation of properties of neighboring Gaussians. The final loss is as follows:
$\mathcal{L_{\mathrm{Gaussians}}}=L_{2}+\lambda_{LPIPS}L_{\textrm{LPIPS}}+\lambda_{SSIM}L_{\textrm{SSIM}}+\lambda_{Sobel}L_{\textrm{Sobel}}+\lambda_{KNN}L_{\textrm{KNN}}.$ (2)
@@
The transparency regularization pushes opacity $o_{i}$ of Gaussians down as following:
@@
$\mathcal{L_{\mathrm{filtering}}}=\mathcal{L_{\mathrm{Gaussians}}}+\lambda_{opacity}L_{\textrm{opacity}}+\lambda_{dice}L_{\textrm{dice}}(S_{\mathrm{GT}},bin(\mathcal{D})||\mathcal{A}).$ (5)
As a result of such training, we remove only Gaussians that could be replaced with the underlying mesh.
@@ 3.3 Merging Gaussians with mesh representation
where $\alpha_{i}[x,y]$ initially calculates based on the opacity $o_{i}^{j}$ and the Gaussian attenuation. We also store the final Gaussians transparency map for each pixel to the alpha map $\mathcal{A}$. To do this, we accumulate transparency at each $[x,y]$ pixel during 3DGS rasterization [31]:
@@ 4 Experiments
In our experiments, we compared HAHA to the state-of-the-art Gaussian methods, namely: GART [38], 3dgs-Avatar [49], and GaussianAvatar [20]. All these methods represent the human body as a set of Gaussians. We used two open datasets to evaluate our approach: SnapshotPeople [7] and X-Humans [53]. From both datasets, we used monocular RGB videos as input to our method. In the following section, we show both qualitative and quantitative results.
@@ 4.1 Implementation details
In our experiments, we set loss weight values as follows. We set $\lambda_{LPIPS}=0.01,$ $\lambda_{SSIM}=0.1,\lambda_{Sobel}=1.0,\lambda_{KNN}=0.01,L_{\mathrm{TV}}$ to $\lambda_{TV}=0.01$ for all stages’ losses i.e. $\mathcal{L_{\mathrm{Gaussians}}}$, $\mathcal{L_{\mathrm{texture}}}$ and $\mathcal{L_{\mathrm{filtering}}}$. We get the best trade-off in quality and number of Gaussians for learnable removing with the following regularization weights: $\lambda_{opacity}=0.001,\lambda_{dice}=0.1$. We trained all avatars using batch size equal to 4 using Adam [33] optimizer. In the first stage of training, we optimize Gaussian parameters for 3000 iterations. Then we optimize texture in the second stage for 2500 iterations. In the last stage, we fine-tuned Gaussians’ color and opacity for 5000 iterations. We set other hyperparameters (such as learning rates) similar to GART [38]. For more details, please refer to the supplementary materials and our source code. [^1] Source code https://github.com/david-svitov/HAHA/
4.2 PeopleSnapshot
Following the previous literature, we provide quantitative metrics for the SnapshotPeople [7] dataset: PSNR, SSIM, and LPIPS [64] (Table 1). PSNR and SSIM measure the fidelity of the signal and structural similarity, respectively, while LPIPS correlates with human perception of the image using neural network features to compare with ground truth. We evaluated metrics on the renderings with a black background as in 3dgs-Avatar [49] experiments to set the background value to zero. For our method, we used test time pose optimization following GART [38] to reduce the impact of SMPL-X fitting inaccuracies.
Figure 5: Novel poses for SnapshotPeople dataset. HAHA reduces the number of artifacts for novel views and body regions unseen during training. At the same time, we use fewer Gaussians for rendering than GART [38].
Table 1: Quantitative metrics for SnapshotPeople [7] dataset. Our method gets metrics on par with state-of-the-art approaches while using much less Gaussians.
male-3-casual female-3-casual
[…]
GART [38] 21.88k 35.93 0.9767 0.0294 19.67k 32.73 0.9672 0.0459
3dgs-Avatar [49] 37.22k 34.28 0.9724 0.0149 53.78k 30.57 0.9581 0.0208
GaussianAvatar [20] 202.73k 33.59 0.9697 0.0243 202.73k 25.94 0.9673 0.0434
HAHA(Ours) 13.60k 31.46 0.9619 0.0277 13.67k 32.53 0.9633 0.0403
male-4-casual female-4-casual
[…]
GART [38] 20.37k 33.17 0.9719 0.0376 21.22k 34.42 0.9706 0.0328
HAHA(Ours) 14.64k 27.08 0.9505 0.0432 11.96k 31.15 0.9589 0.0283
In all experiments for SnapshotPeople we used SMPL provided by AnimNerf [11]. As our method requires a parametric model to have articulated fingers, we converted the provided SMPL to SMPL-X using a converter from the SMPL official repository. Then we fine-tuned the resulting SMPL-X hand’s pose and shape using SMPLify-X[46] to match ground truth frames.
SnapshotPeople evaluation methodology is challenging for our method because we strongly rely on the underlying mesh geometry. Therefore in cases when train and test views and poses are similar, we could face metrics value reduction on the opposite to the methods where rendering does not strongly depend on the mesh surface. Nevertheless, we demonstrate metrics on par with state-of-the-art approaches for this dataset while using almost two times fewer Gaussians (Table 1). In Figure 4 we provide a qualitative comparison of avatar reconstruction for test frames from the PeopleSnapshot dataset. Our method demonstrates on par quality of the avatar using fewer Gaussians (Table 1). Additionally, in some regions that are difficult to represent with Gausssians based on limited input data, our method reduces the number of artifacts (Fig. 5).
The qualitative improvement to the state-of-the-art is noticeable for novel poses and viewpoints (Fig. 5). To demonstrate how our and competitors’ methods handle novel poses, we provide a comparison with reposed results. The use of textured mesh prior not only allows us to reduce the number of Gaussians but also reduces artifacts, especially in areas not sufficiently represented in the training frames. Additional qualitative comparisons with 3dgs-Avatar and visualization for more people from Table 1 can be found in the supplementary materials.
4.3 X-Humans
SnapshotPeople train and test sequences are very similar-looking and do not allow the evaluation of metrics for novel views and poses. To demonstrate the efficiency of HAHA in more challenging scenarios, we measured metrics for RGB video sequences from the X-Humans [53] dataset. X-Humans provides a sequence of frames with rendered 3D scans of a person doing complex movements. The movements are diverse for both training and testing videos, therefore it is a harder task to train on such a dataset. In Table 2 we compare our method with previous state-of-the-art methods. Namely, GART [38] as it gets the best PSNR with the fewer Gaussians, and GaussianAvatar [20] as it uses SMPL-X and can control fingers animation. We provide metrics for both male and female avatars.
Similar to SnapshotPeople we evaluate metrics with a black background and use test time pose optimization for both GART and our approach. HAHA is more robust and gets better metrics for these complex training and testing sequences. Besides, our method requires fewer Gaussians. We also provide qualitative results in Figure 6 demonstrating overall quality and how our approach handles hands animation. Additional qualitative and quantitative comparisons for more people can be found in the supplementary materials.
@@ 4.3 X-Humans
Table 2: Quantitative metrics for X-Humans [53] dataset. The dataset let us evaluate metrics values for novel poses.
@@
Figure 7: Losses ablation study. HAHA gets the best trade-off between image quality and the number of Gaussians with the full set of proposed losses. Also removing Sobel loss could help to additionally reduce the amount of Gaussians but it leads to blurried edges.
We also evaluated the quality of avatars after each stage of training. We provided quantitative metrics in Table 4 while qualitative comparison can be found in Figure 3 (a, b, d). In this experiment, we also ablated the effectiveness of unnecessary Gaussians removal. Naive merging is a baseline method that merges Gaussians with a textured mesh without filtering them out. Using more Gaussians leads to higher PSNR and SSIM, while LPIPS is decreasing. As LPIPS is known for better correlation with human perception, we claim that removing part of the Gaussians leads to better quality.
@@
Gaussians $\downarrow$ Storage (Mb) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Full-Gaussian 37.25k 2.086 28.66 0.9601 0.0572
Textured mesh —— 0.196 27.26 0.9554 0.0539
Naive merging 37.25k 2.282 28.51 0.9599 0.0510
Our trainable merging 12.26k 0.883 28.49 0.9593 0.0501
In Table 4 we demonstrate how the reduction in the number of Gaussians affects the storage space consumption and rendering speed. Using fewer Gaussians with $\times256$ RGB texture lets us use more than 2.3x less memory to store an avatar. Such storage space reduction could be useful for industrial applications when it is necessary to store avatars for millions of users.

### v2

@@ Abstract
We present HAHA - a novel approach for animatable human avatar generation from monocular input videos. The proposed method relies on learning the trade-off between the use of Gaussian splatting and a textured mesh for efficient and high fidelity rendering. We demonstrate its efficiency to animate and render full-body human avatars controlled via the SMPL-X parametric model. Our model learns to apply Gaussian splatting only in areas of the SMPL-X mesh where it is necessary, like hair and out-of-mesh clothing. This results in a minimal number of Gaussians being used to represent the full avatar and reduced rendering artifacts. This allows us to handle the animation of small body parts, such as fingers, that are traditionally disregarded. We demonstrate the effectiveness of our approach on two open datasets: SnapshotPeople and X-Humans. Our method demonstrates on par reconstruction quality to the state-of-the-art on SnapshotPeople, while using less than a third of Gaussians. HAHA outperforms previous state-of-the-art on novel poses from X-Humans both quantitatively and qualitatively.
Figure 1: Optimizing the number of Gaussians. HAHA jointly optimizes a Gaussian splatting model with a textured mesh to improve the photometric quality of the avatars. The method filters out superfluous Gaussians in a learnable, unsupervised manner. As a result, we can more efficiently and better animate highly articulated parts of a body.
@@ 1 Introduction
To date, several options are available to generate human avatars for what concerns input data. To get the best quality many methods rely on multi-view data [9, 11, 65, 25, 18]. However, complex acquisition systems such as a multi-camera capture setup [23] or a 3D scanner [4] are required to collect such data. On the other hand, some methods [55, 19, 51, 8] use a single image of a person as input, which overcomplicates the task with the necessity to restore unobserved regions of the body. Eventually, the most convenient way is to generate avatars from monocular videos. Using a monocular video provides a trade-off between the complexity of obtaining input data and the quality of the avatar.
@@
  • We develop an unsupervised method for significantly reducing the amount of Gaussians in the scene through the use of textured mesh;
  • We demonstrate that our method can efficiently handle the animation of hands and other highly articulated parts without the need for any additional engineering.
@@ 2 Related Work
Several previous works [67, 53] solve the task of generating human avatars with articulated finders using multi-view data. AvatarReX [67] uses a separate parametric model to process hands. As input, they accept a multi-view video of a person. X-Avatar [53] uses a part-specific deformer network to handle the hands. As input X-Avatar gets 3D scans or RGB-D video with depth information. Both methods reconstruct an avatar as a textured mesh that in general leads to more blurred results than 3DGS. In contrast to the previous works, we use only monocular RGB data.
@@
In this work, we research the new task of merging a novel Gaussian-based approach with a classical RGB texture-based. This allows us to reduce the number of Gaussians and, therefore, reduce memory requirements to store an avatar. Utilizing textured mesh where possible helps us reduce the number of artifacts connected with redundant Gaussians while remaining Gaussians represent out-of-the-mesh elements of avatars. Thus, we leverage the pros from both representations.
@@ 3.2 Monocular Avatar Training
Figure 3: Stages of training. a) SMPL-X with optimizable RGB texture fitted on input video frames. b) 3DGS trained as described in Section 3.1. c) All unnecessary Gaussians are deleted (Section 3.3) to merge this step with (a) and get (d).
First stage: Gaussian avatar training. In the first stage (Fig. 2 (a)), we train the 3DGS representation of an avatar by optimizing only local Gaussians transformations $\mu_{i}^{j},r_{i}^{j},s_{i}^{j}$ and color $c_{i}^{j}$. Opacity $o_{i}^{j}$ is fixed to $1$ during this stage as we keep all Gaussians not transparent to efficiently back-propagate image space losses to the SMPL-X parameters. Thus, we force the model to optimize the pose and shape of the underlying mesh rather than deleting Gaussians. We use randomly colored backgrounds in this stage to prevent Gaussians from learning background color.
To optimize Gaussians we use several image space losses as $L_{2}$ loss, $L_{\textrm{LPIPS}}$ perceptual loss [63], $L_{\textrm{SSIM}}$ structure similarity loss, and $L_{\textrm{Sobel}}$ loss to get sharper edges. To calculate $L_{\textrm{Sobel}}$ loss we measure $L_{2}$ between results of applying the Sobel operator [30] to rendered and ground truth images. In other words, we calculate the distance between discrete derivatives of images to account for high-frequency details. We follow [38] and apply $L_{\textrm{KNN}}$, a KNN-based regularization to get smoother results with fewer artifacts. In KNN-regularization we minimize the standard deviation of properties of neighboring Gaussians. The final loss is as follows:
$\mathcal{L_{\mathrm{Gaussian}}}=L_{2}+\lambda_{LPIPS}L_{\textrm{LPIPS}}+\lambda_{SSIM}L_{\textrm{SSIM}}+\lambda_{Sobel}L_{\textrm{Sobel}}+\lambda_{KNN}L_{\textrm{KNN}}.$ (2)
@@
The transparency regularization pushes opacity $o_{i}$ of Gaussians down as follows:
@@
$\mathcal{L_{\mathrm{filtering}}}=\mathcal{L_{\mathrm{Gaussian}}}+\lambda_{opacity}L_{\textrm{opacity}}+\lambda_{dice}L_{\textrm{dice}}(S_{\mathrm{GT}},bin(\mathcal{D})||\mathcal{A}).$ (5)
As a result of such training, we remove only Gaussians that could be replaced with the underlying mesh. Finally, in the inference stage, we utilize preserved Gaussians and trained texture to render an avatar driven by SMPL-X pose parameters.
@@ 3.3 Merging Gaussians with Mesh Representation
where $\alpha_{i}[x,y]$ initially calculates based on the opacity $o_{i}^{j}$ and the Gaussian attenuation (For more details, please refer to the supplementary materials). We also store the final Gaussians transparency map for each pixel to the alpha map $\mathcal{A}$. To do this, we accumulate transparency at each $[x,y]$ pixel during 3DGS rasterization [31]:
@@ 4 Experiments
In our experiments, we compared HAHA to the state-of-the-art Gaussian methods, namely: GART [38], 3DGS-Avatar [49], and GaussianAvatar [20]. All these methods represent the human body as a set of Gaussians. We used two open datasets to evaluate our approach: X-Humans [53] and SnapshotPeople [7]. From both datasets, we used monocular RGB videos as input to our method. In the following section, we show both qualitative and quantitative results.
@@ 4.1 Implementation Details
In our experiments, we set loss weight values as follows. We set $\lambda_{LPIPS}=0.01,\lambda_{SSIM}=0.1,\lambda_{Sobel}=1.0,\lambda_{KNN}=0.01,L_{\mathrm{TV}}$ to $\lambda_{TV}=0.01$ for all stages’ losses i.e. $\mathcal{L_{\mathrm{Gaussian}}}$, $\mathcal{L_{\mathrm{texture}}}$ and $\mathcal{L_{\mathrm{filtering}}}$. We get the best trade-off in quality and number of Gaussians for learnable removing with the following regularization weights: $\lambda_{opacity}=0.001,\lambda_{dice}=0.1$. We trained all avatars using batch size equal to 4 using Adam [33] optimizer. In the first stage of training, we optimize Gaussian parameters for 3000 iterations. Then we optimize texture in the second stage for 2500 iterations. In the last stage, we fine-tuned Gaussians’ color and opacity for 5000 iterations. We set other hyperparameters (such as learning rates) similar to GART [38]. For more details, please refer to supplementary materials.
@@
Table 1: Quantitative metrics for X-Humans [53] dataset. The dataset lets one evaluate metrics values for novel poses.
@@
00018 (male) 00027 (female)
[…]
3DGS-Avatar [49] 26.78k 28.71 0.9521 0.0580 36.82k 26.84 0.9477 0.0445
GART [38] 50.47k 30.98 0.9595 0.0683 47.18k 26.56 0.9449 0.0595
GaussianAvatar [20] 191.58k 29.92 0.9588 0.0744 191.58k 25.69 0.9481 0.0543
HAHA(Ours) 18.57k 31.10 0.9630 0.0579 15.50k 27.26 0.9513 0.0473
Table 2: Quantitative metrics for SnapshotPeople [7] dataset. Our method gets metrics on par with state-of-the-art approaches while using much fewer Gaussians.
female-3-casual male-3-casual
[…]
3DGS-Avatar [49] 53.78k 30.57 0.9581 0.0208 37.22k 34.28 0.9724 0.0149
GART [38] 19.67k 32.73 0.9672 0.0459 21.88k 35.93 0.9767 0.0294
GaussianAvatar [20] 202.73k 25.94 0.9673 0.0434 202.73k 33.59 0.9697 0.0243
HAHA(Ours) 13.67k 32.53 0.9633 0.0403 13.60k 31.46 0.9619 0.0277
4.2 X-Humans
We report the following metrics: PSNR, SSIM, and LPIPS [64] (Table 1). PSNR and SSIM measure the fidelity of the signal and structural similarity, respectively, while LPIPS correlates with human perception of the image using neural network features to compare with ground truth. We evaluated metrics on the renderings with a black background as in 3DGS-Avatar [49] experiments to set the background value to zero. During inference, we used test time pose optimization following GART [38] to reduce the impact of SMPL-X fitting inaccuracies.
X-Humans [53] dataset provides a sequence of frames with rendered 3D scans of a person doing complex movements. The movements are diverse for both training and testing videos, therefore it is a challenging task to train on such a dataset. In Table 1 we compare our method with previous state-of-the-art methods: GART [38], 3DGS-Avatar [49], and GaussianAvatar [20]. The last one, similar to us, uses SMPL-X and can control fingers animation so we can compare the animation of hands. We provide metrics for both male and female avatars.
HAHA is more robust and gets better metrics for these complex training and testing sequences. Besides, our method requires fewer Gaussians. We also provide qualitative results in Figure 5 demonstrating overall quality and how our approach handles hands animation. Additional visualizations for more people from the dataset can be found in the supplementary materials.
Figure 6: Novel poses for SnapshotPeople dataset. HAHA reduces the number of artifacts for novel views and body regions unseen during training. At the same time, we use fewer Gaussians for rendering.
4.3 PeopleSnapshot
Following the previous literature, we also provide quantitative metrics for the SnapshotPeople [7] dataset (Table 5). However, SnapshotPeople does not allow assess quality for novel views and poses since train and test sequences are very similar-looking.
In all experiments for SnapshotPeople we used SMPL provided by AnimNerf [11]. As our method requires a parametric model to have articulated fingers, we converted the provided SMPL to SMPL-X using a converter from the SMPL official repository. Then we fine-tuned the resulting SMPL-X hand’s pose and shape using SMPLify-X[46] to match ground truth frames. Similar to X-Humans experiments, we evaluate metrics with a black background and use test time pose optimization during inference.
SnapshotPeople evaluation methodology is challenging for our method because we strongly rely on the underlying mesh geometry. Therefore, in cases when train and test views and poses are similar, we could face metrics value reduction on the opposite to the methods where rendering does not strongly depend on the mesh surface. Nevertheless, we demonstrate metrics on par with state-of-the-art approaches for this dataset while using almost two times fewer Gaussians (Table 5). In Figure 4 we provide a qualitative comparison of avatar reconstruction for test frames from the PeopleSnapshot dataset. Our method demonstrates on par quality of the avatar using fewer Gaussians. Additionally, in some regions that are difficult to represent with Gausssians based on limited input data, our method reduces the number of artifacts (Fig. 6).
The qualitative improvement to the state-of-the-art is noticeable for novel poses and viewpoints (Fig. 6). To demonstrate how our and competitors’ methods handle novel poses, we provide a comparison with reposed results. The use of textured mesh prior not only allows us to reduce the number of Gaussians but also reduces artifacts, especially in areas not sufficiently represented in the training frames. Additional visualizations and metrics for more people from the dataset can be found in the supplementary materials.
@@ 4.3 PeopleSnapshot
Figure 7: Losses ablation study. HAHA gets the best trade-off between image quality and the number of Gaussians with the full set of proposed losses. Also removing Sobel loss could help to additionally reduce the amount of Gaussians but it leads to blurred edges.
We also evaluated the quality of avatars after each stage of training. We provided quantitative metrics in Table 4 while qualitative comparison can be found in Figure 3 (a, b, d). In this experiment, we also ablated the effectiveness of unnecessary Gaussians removal. Naive merging is a baseline method that merges Gaussians with a textured mesh without filtering them out. Using more Gaussians leads to higher PSNR and SSIM, while LPIPS decreases. As LPIPS is known for better correlation with human perception, we claim that removing part of the Gaussians leads to better quality.
@@
[…]
(Mb) (Inference) (min:sec)
Textured mesh —— 0.196 $463.58\pm 17.46$ 5:24 27.26 0.9554 0.0539
Full-Gaussian 37.25k 2.086 $240.09\pm 7.65$ 7:58 28.66 0.9601 0.0572
Naive merging 37.25k 2.282 $238.17\pm 7.81$ —— 28.51 0.9599 0.0510
Finetuning 12.26k 0.883 $247.87\pm 4.62$ 12:19 28.49 0.9593 0.0501
In Table 4 we demonstrate how the reduction in the number of Gaussians affects the storage space consumption and rendering speed. Using fewer Gaussians with $\times256$ RGB texture lets us use more than 2.3x less memory to store an avatar. Such storage space reduction could be useful for industrial applications when it is necessary to store avatars for millions of users. At the same time, we demonstrate total convergence speed (25 min 41 sec) on par with other methods: GaussianAvatar (28 min 37 sec), 3DGS-Avatar (26 min 03 sec).